what is the volume of hydrogen gas if there are 1.37 moles at 842 mmHg and 25 degree Celsius


1
Expert's answer
2021-01-18T01:33:06-0500

According to ideal gas law:


"PV=nRT,\\\space\\
V=\frac{nRT}{PV},\\\space\\
V=\frac{1.37\cdot8.314\cdot(25+273)}{842\cdot133.322}=0.03\text{ m}^2."